4. (**) The market demand function is: P= 240-4Qo. For each of the following total cost functions, fill in the table: Note: For submitting this HW, only fill in columns (1) and (2) of the table. Only these two columns will be graded. [1 6-6 points (a) C (Q) = Q2 + 40Q + 150 (b) C(Q) = 40Q + 150 (c) C(Q) 150 Monopolist Monopolistunder underI TCS TPS Monopoly Monopoly under monopoly onaly market
Answer #1

The market demand function: Pb 240-4Qd TR = 240Q-402 MR 240-8Q TC = Q^2 + 400 + 150MC = 2Q +40 MR = MC 240-80-2Q 40 Q=20 By substituting the value of Q in the price equation, P 240-4Q P 240-4*20 P= 160 TC = 4By substituting the value of Q in the price equation, P-240-4Q P = 240-4*25 P 140 C. TC= 150 MC 0 MR MC 240-80 = 0 Q-30 By suMonopolist quantity Monopolist price 20 25 30 160 1A0 120